# Token Interdependency Parsing (Tipping) - Fast and Accurate Log Parsing

### v1

@@ 4.3.3 Summary
To summarize, we conducted speed test experiments on two computing platforms with nine datasets, totaling 15 experiments, as we were unable to run the three largest datasets on our local machine. We found that Tipping was the fastest parser in 12 out of the 15 experiments. In two experiments on the local machine (HDFS and WebShop), Tipping was second to PlIplom but then performed faster on the virtual machine, which has more CPU cores, though the cores themselves are slower. A similar result was observed in the local machine with TrainTicket, where Drain was the fastest. This provides evidence of Tipping’s superior parallelization capability. The second fastest parser was PlIplom, which secured two first places. Drain was the third fastest, with one first place. The notable difference between PlIplom and Drain is that PlIplom tends to be faster than Drain in medium and large datasets, while Drain performs better in small datasets.
@@ 4.4.1 Experiment depiction
In our final experiment, we examine how the parsing output of different parsers affects the downstream task of anomaly detection in software logs. Past work provides somewhat mixed results on whether the accuracy of log parsing impacts the performance of anomaly detection [42]. Nevertheless, in a real-world scenario, the accuracy of the end goals is an important property of analysis pipelines, as ideally, one chooses components based on parsing performance and the end goal’s accuracy.
@@
We tested a combination of supervised and unsupervised algorithms to determine if parsing results impact anomaly detection accuracy. We used the Decision Tree (su-DT), Linear Support-Vector Machine (su-SVM), Logistic Regression (su-LR), Random Forest (su-RF), and Extreme Gradient Boosting Tree (su-XGB) for supervised methods, and K-means clustering (us-KM) and Isolation Forest (us-IF) for unsupervised methods. Additionally, we tested a custom-made unsupervised model called RarityModel (us-RM) [43] that utilizes the negative logarithm to measure the rarity of individual log events. This model proposes that the rarer a log event, the higher the likelihood of an anomaly.
@@ 4.4.2 Experiment results
Hadoop results are presented in Tables XXII and XXIII. PlIpLom achieves the best results in terms of both $F_{1}$ score and AUC-ROC, followed by Tipping, while the last place alternates between Drain and Brain. The difference between the winner, PlIpLom, and the last parser in the AUC-ROC average is 0.016 (0.612 - 0.596) and in $F_{1}$ score 0.024 (0.337 - 0.313). PlIpLom’s top performance stems from its results with the SVM model, where it significantly outperforms the last parser with an AUC-ROC of 0.825 compared to 0.750 and an $F_{1}$ score of 0.574 compared to 0.377. The differences between PlIpLom and other parsers are negligible with all other machine learning models.
TABLE XXII: AUC-ROC of anomaly detection on Hadoop
@@
TABLE XXIII: $F_{1}$ score of anomaly detection on Hadoop
@@
Tables XXIV and XXV demonstrate the HDFS results, showing no meaningful differences between parsers’ average $F_{1}$ scores or AUC-ROCs. However, when examining individual machine learning models and parsers pairs, differences in terms of $F_{1}$ score are observed (see Table XXV). Brain shows slightly weaker performance with LR and RF compared to the rest but compensates with better performance using the IF model. The notable differences between machine learning algorithms suggest that the setup should allow for detecting differences between parsers if there are any.
TABLE XXIV: AUC-ROC of anomaly detection on HDFS
@@
TABLE XXV: $F_{1}$ score of anomaly detection on HDFS
@@
BGL results are in Tables XXVI and XXVII. Tipping is the winner while the margin to the last parser in AUC-ROC is 0.068 (0.696 - 0.628) and in $F_{1}$ score 0.104 (0.329 - 0.225). Tipping is superior with all supervised algorithms and IF from the unsupervised algorithms. With unsupervised algorithms, Brain is better than Tipping with RM and is on par with KM.
TABLE XXVI: AUC-ROC of anomaly detection on BGL
@@
TABLE XXVII: $F_{1}$ score of anomaly detection on BGL
@@ 4.4.3 Summary
Past work [42] has proposed that the number of templates a parser generates is negatively correlated with the anomaly detection capability of the parser. We checked the number of events the four parsers produced, as shown in Table XXVIII, and found mixed results. In HDFS, Drain produces 43% more events than Tipping, but there is no difference in anomaly detection capability in that dataset. In Hadoop, PlIplom was the best parser, but it did not produce the smallest amount of events. This theory only holds in the BGL dataset, where Tipping produced the smallest number of events and was the best parser.
Table XXVIII also compares the ”correct” number of templates from two sources produced by humans on the full dataset [1, 44]. Event numbers from the sources do not match each other, indicating that it is hard to arrive at the same result, even when we agree on the concept of templates and parameters on a higher level. Using those numbers suggests that all parsers are in the right ballpark regarding HDFS and Hadoop. However, regarding BGL, many parsers seem to produce too many templates. Thus, this would bring forward a new hypothesis that the number of templates decreases the anomaly detection accuracy only if the number deviates too much from the ”correct” number of templates.
TABLE XXVIII: Number of events produced
Dataset Brain Drain PlIpLom Tip [44] [1]

### v2

@@ 4.4.3 Summary
To summarize, we conducted speed test experiments on two computing platforms with nine datasets, totaling 15 experiments, as we could not run the three most enormous datasets on our local machine. We found that Tipping was the fastest parser in 12 out of the 15 experiments. In two experiments on the local machine (HDFS and WebShop), Tipping was second to PlIplom but then performed faster on the virtual machine, which has more CPU cores, though the cores are slower. A similar result was observed in the local machine with TrainTicket, where Drain was the fastest. This provides evidence of Tipping’s superior parallelization capability. The second fastest parser was PlIplom, which secured two first places. Drain was the third fastest, with one first place. The notable difference between PlIplom and Drain is that PlIplom tends to be faster than Drain in medium and large datasets, while Drain performs better in small datasets.
@@ 4.5.1 Experiment depiction
In our final experiment, we examine how the parsing output of different parsers affects the downstream task of anomaly detection in software logs. We used the LogLead tool described in Section 4.4 to carry out the experiments.
Past work provides somewhat mixed results on whether the accuracy of log parsing impacts the performance of anomaly detection [parsereffect]. Nevertheless, in a real-world scenario, the accuracy of the end goals is an important property of analysis pipelines, as ideally, one chooses components based on parsing performance and the end goal’s accuracy.
@@
We tested a combination of supervised and unsupervised algorithms to determine if parsing results impact anomaly detection accuracy. We used the Decision Tree (su-DT), Linear Support-Vector Machine (su-SVM), Logistic Regression (su-LR), Random Forest (su-RF), and Extreme Gradient Boosting Tree (su-XGB) for supervised methods, and K-means clustering (us-KM) and Isolation Forest (us-IF) for unsupervised methods. Additionally, we tested a custom-made unsupervised model called RarityModel (us-RM) [nyyssola2023efficiency] that utilizes the negative logarithm to measure the rarity of individual log events. This model proposes that the rarer a log event, the higher the likelihood of an anomaly.
@@ 4.5.2 Experiment results
Hadoop results are presented in Tables XII and XIII. PlIpLom achieves the best results in terms of both $F_{1}$ score and AUC-ROC, followed by Tipping, while the last place alternates between Drain and Brain. The difference between the winner, PlIpLom, and the last parser in the AUC-ROC average is 0.016 (0.612 - 0.596) and in $F_{1}$ score 0.024 (0.337 - 0.313). PlIpLom’s top performance stems from its results with the SVM model, where it significantly outperforms the last parser with an AUC-ROC of 0.825 compared to 0.750 and an $F_{1}$ score of 0.574 compared to 0.377. The differences between PlIpLom and other parsers are negligible with all other machine learning models.
TABLE XII: AUC-ROC of anomaly detection on Hadoop
@@
TABLE XIII: $F_{1}$ score of anomaly detection on Hadoop
@@
Tables XIV and XV demonstrate the HDFS results, showing no meaningful differences between parsers’ average $F_{1}$ scores or AUC-ROCs. However, when examining individual machine learning models and parsers pairs, differences in terms of $F_{1}$ score are observed (see Table XV). Brain shows slightly weaker performance with LR and RF compared to the rest but compensates with better performance using the IF model. The notable differences between machine learning algorithms suggest that the setup should allow for detecting differences between parsers if there are any.
TABLE XIV: AUC-ROC of anomaly detection on HDFS
@@
TABLE XV: $F_{1}$ score of anomaly detection on HDFS
@@
BGL results are in Tables XVI and XVII. Tipping is the winner while the margin to the last parser in AUC-ROC is 0.068 (0.696 - 0.628) and in $F_{1}$ score 0.104 (0.329 - 0.225). Tipping is superior with all supervised algorithms and IF from the unsupervised algorithms. With unsupervised algorithms, Brain is better than Tipping with RM and is on par with KM.
TABLE XVI: AUC-ROC of anomaly detection on BGL
@@
TABLE XVII: $F_{1}$ score of anomaly detection on BGL
@@ 4.5.3 Summary
Past work [parsereffect] has proposed that the number of templates a parser generates is negatively correlated with the anomaly detection capability of the parser. We checked the number of events the four parsers produced, as shown in Table XVIII, and found mixed results. In HDFS, Drain produces 43% more events than Tipping, but there is no difference in anomaly detection capability in that dataset. In Hadoop, PlIplom was the best parser, but it did not produce the smallest amount of events. This theory only holds in the BGL dataset, where Tipping produced the smallest number of events and was the best parser.
Table XVIII also compares the ”correct” number of templates from two sources produced by humans on the full dataset [logpm, landauer2023critical]. Event numbers from the sources do not match each other, indicating that it is hard to arrive at the same result, even when we agree on the concept of templates and parameters on a higher level. Using those numbers suggests that all parsers are in the right ballpark regarding HDFS and Hadoop. However, regarding BGL, many parsers seem to produce too many templates. Thus, this would bring forward a new hypothesis that the number of templates decreases the anomaly detection accuracy only if the number deviates too much from the ”correct” number of templates.
TABLE XVIII: Number of events produced
Dataset Brain Drain PlIpLom Tip [landauer2023critical] [logpm]
@@ 5.1 Internal Validity
Although Tipping’s and the benchmark’s implementations undergo rigorous testing, the outcomes may inadvertently indicate a bug which may invalidate evaluations.